# Stochastic Sampling of Operator Growth Dynamics

### v1

@@ I.2 Swap
I.3 Interact
Figure S3: Illustration of the interact move. Two $\mathcal{L}_{ZZ}$ Liouvillians sharing only one site may “interact” to give a pair of $\mathcal{L}_{Z}$ Liouvillians situated on next to nearest neighbour sites.
The add-drop move only allows Liouvillians to exist on a site in pairs. This can however miss configurations with an odd number of bond Liouvillians on a site. To enable this we introduce the interact move, as illustrated in Fig. S3. It is to be noted that for $d>1$, the two interacting bonds can be along different spatial axes. In such cases, two different $\mathcal{L}_{ZZ}$ pair configurations can map to the same $\mathcal{L}_{Z}$ pair configuration. This needs to be accounted for in detailed balance.
@@ I.3 Interact
    If the Liouvillian on both sides of the time slice are of type $\mathcal{L}_{ZZ}$ and share exactly one site, check if the shared site has a $Z$ or $I$ operator on it. If so, replace the $\mathcal{L}_{ZZ}$ Liouvillians with $\mathcal{L}_{Z}$ Liouvillians as in Fig. S3.
@@
    If the Liouvillian slots on both sides host $\mathcal{L}_{Z}$ Liouvillians placed at next to nearest neighbour sites, check if their common nearest neighbouring site has a $Z$ or $I$ operator. If this is the case, replace the $\mathcal{L}_{Z}$ Liouvillians with $\mathcal{L}_{ZZ}$ Liouvillians as in Fig. S3.
@@
$\frac{T_{i\rightarrow f}A_{i\rightarrow f}}{T_{f\rightarrow i}A_{f\rightarrow i}}=\frac{P_{f}}{P_{i}}=\frac{(h^{Z})^{2}}{J^{2}}$ (S10)
@@
$T_{i\rightarrow f}=\frac{1}{2n_{\max}+1}\,,\,T_{f\rightarrow i}=\frac{1}{(2n_{\max}+1)\alpha}\,,$ (S11)
@@
where $\alpha=2$ if the $\mathcal{L}_{ZZ}$ Liouvillians in configuration $i$ are along different spatial directions (otherwise, $\alpha=1$). This is because, proposing a $\mathcal{L}_{ZZ}\to\mathcal{L}_{Z}$ interact move amounts to choosing a time slice. Meanwhile, to specify a $\mathcal{L}_{Z}\to\mathcal{L}_{ZZ}$ interact move, we need choose a time slice, and then pick a common neighboring site among the $\alpha$ choices. Therefore, the following acceptance rates fulfill detailed balance:
$A_{i\rightarrow f}=\text{min}\left(1,\frac{(h^{Z})^{2}}{\alpha J^{2}}\right)\,,\,A_{f\rightarrow i}=\text{min}\left(1,\frac{\alpha J^{2}}{(h^{Z})^{2}}\right)\,.$ (S12)
II The $X$ Operator
@@
Figure S4: (a) $r_{n}$ vs $n$ for the $X$ operator with $J=h^{Z}=h^{X}=1$ in 1D. The inset focuses in on the first 30 moments, comparing our MCMC with exact numerics. (b) $\Delta r_{n}$ and $\log(n)\Delta r_{n}$ vs $n$ for 1D. (c) Same as (a) for 2D. The inset has the first 12 moments. (d) $\Delta r_{n}$ vs $n$ for 2D.
@@
The statistical nature of infinite temperature operator growth dynamics in non-integrable systems like the mixed field QIM implies that local operators should typically give rise to the same generic behavior at a large enough order $n$. Based on the Operator Growth Hypothesis [21], one expects local operators to asymptotically grow linearly with the same slope. To check this, we compute the same plots as Fig. 3 but for the $X$ operator in Fig. S4. We see that for 1D $\log(n)\Delta r_{n}$ vs $n$ again gives us a constant curve $\log(n)\Delta r_{n}\simeq 2.2(1)J$, the same as the $Z$ operator in the main text. Similarly, in 2D we again get a plot consistent with linear growth with slope $\simeq 1.26(4)J$. Another point to note is the non-universal nature of the intermediate plateau in the 2D $\Delta r_{n}$ vs $n$ plot. We see that here the plateau is much narrower than the one for the $Z$ operator as shown in the main text.
III Additional move for the TFIM Limit
@@
Figure S5: An example of a configuration in the $n=2$ sector that requires the interact move to be generated. Temporal slices of the 2D lattice are indexed by k. The first temporal slice is identical to the last because of the fixed boundary conditions. The $\mathcal{L}_{ZZ}$ Liouvillians, shown as rectangular boxes, transform local Pauli operators between slices.
In the $h^{Z}=0$ limit, the interact move breaks down. This is not a problem in 1D as configurations with an odd number of identical $\mathcal{L}_{ZZ}$ Liouvillians also require the presence of $\mathcal{L}_{Z}$ Liouvillians and thus do not exist in this limit. For $d>1$ it is possible for configurations to have an odd number of identical $\mathcal{L}_{ZZ}$ Liouvillians without $\mathcal{L}_{Z}$ Liouvillians. Fig. S5 shows a simple example with $\mathcal{L}_{ZZ}$ Liouvillians looping around a plaquette, satisfying the boundary conditions. Note that each site hosts just one $\mathcal{L}_{ZZ}$ Liouvillian of a kind. One can generate larger loops by removing $\mathcal{L}_{ZZ}$ pairs on the common edges of overlapping smaller loops using add-drop.
@@
Figure S6: Section of two configurations connected by the rotate move on a 2D lattice. k corresponds to the temporal index. Note that both configurations have the same temporal boundaries, implying the local nature of the move.
To access this part of the configuration space in absence of interact we propose the rotate move. Fig. S6 shows that it is possible to rotate two parallel $\mathcal{L}_{ZZ}$ Liouvillians sharing a plaquette to get another pair of parallel $\mathcal{L}_{ZZ}$ Liouvillians on the same plaquette while maintaining the same boundaries. Detailed balance for this move should be trivial as the involved configurations have the same weight and the Metropolis acceptance rate is $100\%$. It can be seen that a combination of add-drop and rotate can generate the configuration in Fig. S5. We benchmark this move against exact numerics in Fig. S7.
@@
Figure S7: Comparing our MCMC equipped with and without the rotate move against exact computations for the 2D TFIM with the same boundary operator as in Fig. S5.

### v2

@@ A.2 Swap
A.3 Interact
Figure 8: Illustration of the interact move. Two $\mathcal{L}_{ZZ}$ Liouvillians sharing only one site may “interact” to give a pair of $\mathcal{L}_{Z}$ Liouvillians situated on next to nearest neighbour sites.
The add-drop move only allows Liouvillians to exist on a site in pairs. This can however miss configurations with an odd number of bond Liouvillians on a site. To enable this we introduce the interact move, as illustrated in Fig. 8. It is to be noted that for $d>1$, the two interacting bonds can be along different spatial axes. In such cases, two different $\mathcal{L}_{ZZ}$ pair configurations can map to the same $\mathcal{L}_{Z}$ pair configuration. This needs to be accounted for in detailed balance.
@@ A.3 Interact
    If the Liouvillian on both sides of the time slice are of type $\mathcal{L}_{ZZ}$ and share exactly one site, check if the shared site has a $Z$ or $I$ operator on it. If so, replace the $\mathcal{L}_{ZZ}$ Liouvillians with $\mathcal{L}_{Z}$ Liouvillians as in Fig. 8.
@@
    If the Liouvillian slots on both sides host $\mathcal{L}_{Z}$ Liouvillians placed at next to nearest neighbour sites, check if their common nearest neighbouring site has a $Z$ or $I$ operator. If this is the case, replace the $\mathcal{L}_{Z}$ Liouvillians with $\mathcal{L}_{ZZ}$ Liouvillians as in Fig. 8.
@@
$\frac{T_{i\rightarrow f}A_{i\rightarrow f}}{T_{f\rightarrow i}A_{f\rightarrow i}}=\frac{P_{f}}{P_{i}}=\frac{(h^{Z})^{2}}{J^{2}}$ (19)
@@
$T_{i\rightarrow f}=\frac{1}{2n_{\max}+1}\,,\,T_{f\rightarrow i}=\frac{1}{(2n_{\max}+1)\alpha}\,,$ (20)
@@
where $\alpha=2$ if the $\mathcal{L}_{ZZ}$ Liouvillians in configuration $i$ are along different spatial directions (otherwise, $\alpha=1$). This is because, proposing a $\mathcal{L}_{ZZ}\to\mathcal{L}_{Z}$ interact move amounts to choosing a time slice. Meanwhile, to specify a $\mathcal{L}_{Z}\to\mathcal{L}_{ZZ}$ interact move, we need to choose a time slice, and then pick a common neighboring site among the $\alpha$ choices. Therefore, the following acceptance rates fulfill detailed balance:
$A_{i\rightarrow f}=\text{min}\left(1,\frac{(h^{Z})^{2}}{\alpha J^{2}}\right)\,,\,A_{f\rightarrow i}=\text{min}\left(1,\frac{\alpha J^{2}}{(h^{Z})^{2}}\right)\,.$ (21)
Appendix B $X$ Correlation Function
@@
Figure 9: (a) $r_{n}$ vs $n$ for the $X$ operator with $J=h^{Z}=h^{X}=1$ in 1D. The inset focuses in on the first 30 moments, comparing our MCMC with exact numerics. (b) $\Delta r_{n}$ and $\log(n)\Delta r_{n}$ vs $n$ for 1D. (c) Same as (a) for 2D. The inset has the first 12 moments. (d) $\Delta r_{n}$ vs $n$ for 2D.
@@
The statistical nature of infinite temperature operator growth dynamics in non-integrable systems like the mixed field QIM implies that local operators should typically give rise to the same generic behaviour at a large enough order $n$. Based on the Operator Growth Hypothesis [21], one expects local operators to asymptotically grow linearly with the same slope. To check this, we compute the same plots as Fig. 3 in the main text but for the $X$ operator in Fig. 9. We see that for 1D $\log(2n)\Delta r_{n}$ vs $n$ again gives us a constant curve $\log(2n)\Delta r_{n}\simeq 2.2(1)J$, the same as the $Z$ operator in the main text. Similarly, in 2D we again get a plot consistent with linear growth with slope $\simeq 1.26(4)J$. Another point to note is the non-universal nature of the intermediate plateau in the 2D $\Delta r_{n}$ vs $n$ plot. We see that here the plateau is much narrower than the one for the $Z$ operator as shown in the main text.
@@
Figure 10: An example of a configuration in the $n=2$ sector that requires the interact move to be generated. Temporal slices of the 2D lattice are indexed by k. The first temporal slice is identical to the last because of the fixed boundary conditions. The $\mathcal{L}_{ZZ}$ Liouvillians, shown as rectangular boxes, transform local Pauli operators between slices.
Appendix C Additional move for the TFIM Limit
@@
Figure 11: Section of two configurations connected by the rotate move on a 2D lattice. k corresponds to the temporal index. Note that both configurations have the same temporal boundaries, implying the local nature of the move.
In the $h^{Z}=0$ limit, the interact move breaks down. This is not a problem in 1D as configurations with an odd number of identical $\mathcal{L}_{ZZ}$ Liouvillians also require the presence of $\mathcal{L}_{Z}$ Liouvillians and thus do not exist in this limit. For $d>1$ it is possible for configurations to have an odd number of identical $\mathcal{L}_{ZZ}$ Liouvillians without $\mathcal{L}_{Z}$ Liouvillians. Fig. 10 shows a simple example with $\mathcal{L}_{ZZ}$ Liouvillians looping around a plaquette, satisfying the boundary conditions. Note that each site hosts just one $\mathcal{L}_{ZZ}$ Liouvillian of a kind. One can generate larger loops by removing $\mathcal{L}_{ZZ}$ pairs on the common edges of overlapping smaller loops using add-drop.
@@
Figure 12: Comparing our MCMC equipped with and without the rotate move against exact computations for the 2D TFIM with the same boundary operator as in Fig. 10.
To access this part of the configuration space in absence of interact we propose the rotate move. Fig. 11 shows that it is possible to rotate two parallel $\mathcal{L}_{ZZ}$ Liouvillians sharing a plaquette to get another pair of parallel $\mathcal{L}_{ZZ}$ Liouvillians on the same plaquette while maintaining the same boundaries. Detailed balance for this move should be trivial as the involved configurations have the same weight and thus the Metropolis acceptance rate is $100\%$. It can be seen that a combination of add-drop and rotate can generate the configuration in Fig. 10. We benchmark this move against exact numerics in Fig. 12.